# Critical modular lattices in the Gaussian core model

### v1

@@ 5.3. The Barnes-Wall lattice
(41) $\Theta_{\operatorname{BW}_{16}}(\tau)=E^{(1)}_{8}(\tau)+\frac{480}{17}E^{(2)}_{8}(\tau)-\frac{480}{17}C_{8}(\tau).$
To find an upper bound for the coefficients of $\Theta_{\operatorname{BW}_{16}}$, we will find upper bounds for the coefficients of $E^{(1)}_{8}$, $E^{(2)}_{8}$ and $C_{8}$.
@@
(42) $|b^{(1)}_{m}|,|b^{(2)}_{m}|\leq\left|-\frac{16}{B_{8}}\zeta(7)m^{7}\right|\leq 485m^{7}.$
For $C_{8}(\tau)=\sum_{m=0}^{\infty}c_{m}q^{m}$ we use estimates on the coefficients from [17] as collected in Section 4.2.1. By combining (34) with the estimate $d(m)\leq 2\sqrt{m}$ we obtain for $C_{8}$ specifically that
(43) $\begin{split}|c_{m}|\leq 7.6\cdot 10^{6}\cdot m^{4}.\end{split}$
Combining (41), (42) and (43) we get
$\sum_{m=9}^{\infty}|a_{m}|(6m)^{2}e^{-6m}\leq 5.2\cdot 10^{5}\sum_{m=9}^{\infty}m^{9}e^{-6m}+7.8\cdot 10^{9}\sum_{m=9}^{\infty}m^{6}e^{-6m}.$
With Lemma 5.2 we find
$\sum_{m=9}^{\infty}m^{9}e^{-6m}\leq 1.7\cdot 10^{-15}\quad\text{and}\quad\sum_{m=9}^{\infty}m^{6}e^{-6m}\leq 2.3\cdot 10^{-18}.$
Putting everything together, we finally have
$\left|\frac{1}{288}\sum_{m=9}^{\infty}a_{m}6m(6m-(n/2+1))e^{-6m}\right|\leq 1.2\cdot 10^{-10}<0.0016.$
Thus, all eigenvalues of the Hessian of the energy function $\mathcal{E}(3,L)$ are negative at $L=\operatorname{BW}_{16}$, and we have proven that the Barnes-Wall lattice is a local maximizer for $\alpha=3$. ∎
5.4. Extremal modular lattices which are not critical
@@
Note that, up to isometry, there are two more such lattices. For the first such lattice, the first entry in the above-mentioned file, call it $L$ for the moment, all shells form spherical $3$-designs; to test this one can use a design test proposed by Goethals and Seidel [13]. Specifically one can check that there are no $\operatorname{O}(L)$-invariant harmonic polynomials of degrees $1,\ldots,3$ for the isometry group $\operatorname{O}(L)$ of $L$; this in turn can be checked by computing the harmonic Molien-series of $L$ which starts $1+u^{4}+2u^{6}+\ldots$, we computed this using the computer algebra system MAGMA [18]. Then Theorem 3.12 in [13] asserts that all $\operatorname{O}(L)$-orbits are spherical $3$-designs, and so all shells of $L$ are, as they clearly are unions of orbits of $\operatorname{O}(L)$. So this lattice is a critical point.
@@
This follows the same lines as the computations for the Coxeter-Todd and Barnes-Wall lattice as above. Note that $\Theta_{L_{12}}$ is a modular form of weight 6 for $\Gamma_{0}(2)$. The space $\mathcal{M}_{6}(\Gamma_{0}(2))$ contains no cusp forms and is therefore equal to its two-dimensional subspace of Eisenstein series $\operatorname{Eis}_{6}(\Gamma_{0}(2))$.
@@
5.5. Extremal modular lattices without $4$-designs
We finish the investigation of the local behavior of extremal $2$ and $3$-modular lattices with examples of such lattices where all shells form at least $3$-designs, but not necessarily $4$-designs.
Without proof, we mention that the lowest dimensional example of a lattice which is a local maximum for some Gaussian potential can be found among extremal $2$-modular lattices in dimension $8$, of which there exists a unique isometry-class. For a proof of this, similar to computations in this article, we refer to the Master’s thesis of Joharian [15].
We specifically investigate an extremal $3$-modular lattice $L_{16}$ of dimension $16$, given explicitly as the third entry in “3_dim16.dat” in [22]. The reason we chose this as an example over the $8$-dimensional $2$-modular lattice mentioned above is that here the full machinery of Theorem 3.3 is necessary, making this a more instructive computation.
The theta series of $L_{16}$ is
(50) $\Theta_{L_{16}}(\tau)=1+720q^{2}+13440q^{3}+97200q^{4}+O(q^{5}),$
Proposition 5.5.
For $\alpha=3$ the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(3,L_{16})$ are strictly negative. In particular: $L_{16}$ is a local maximum for $\mathcal{E}(3,L)$ on the manifold of lattices with point density $3^{4}$.
Proof.
To prove this we continue as follows. We firstly bound the coefficients of $\Theta_{L_{16}}$ and then use Theorem 3.3 to estimate the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,L_{16})$ at $L_{16}$, rigorously so for the specific energy parameter $\alpha=3$. For this it turns out that we need the eigenvalues of the auxiliary forms $Q_{4}$ and $Q_{6}$ associated to $L_{16}$, requiring further estimating coefficients of cups forms in $\mathcal{S}_{12}(\Gamma_{0}(3))$.
5.5.1. Bounding the coefficients of $\Theta_{L_{16}}$
Then
$|b^{(1)}|,|b^{(2)}_{m}|\leq|b_{m}|\leq 485m^{7}$
To bound the coefficients of $C_{8}(\tau)$ we again use the estimates coming from [3] as collected in Section 4.2.2. The dimensions of $\mathcal{S}_{8}^{+}(\Gamma_{0}(3))$ and $\mathcal{S}_{8}^{-}(\Gamma_{0}(3))$ are given by
$d_{8}^{+}$ $=\left[\frac{8}{6}\right]=1$
$d_{8}^{-}$ $=\left[\frac{8}{3}\right]-\left[\frac{8}{6}\right]-1=0.$
Since the dimension of $\mathcal{S}_{8}^{-}(\Gamma_{0}(3))$ is zero, we have $C_{8}(\tau)\in\mathcal{S}_{8}^{+}(\Gamma_{0}(3))$.
Using Theorem 1.1 in [3] and $d(m)\leq 2\sqrt{m}$ we find
$|c_{m}|$ $\leq\sqrt{\log(8)}\left(76.585\sqrt{|c_{1}|^{2}}+\frac{e^{39.62}(114.2)^{8}}{6^{3.5}}|c_{1}|e^{-21.821}\right)2m^{4}$
$\leq 8.5\cdot 10^{21}\cdot m^{4}.$
(51) $|a_{m}|\leq 3325\cdot m^{7}+5\cdot 10^{22}\cdot m^{4}.$
5.5.2. The auxiliary forms $Q_{4}$ and $Q_{6}$
$Q_{2m}=\sum_{x\in L_{16}(2m)}H[x]^{2}$
specifically for $m=1,2,3$.
$C^{(3)}_{12}(\tau)=\sum_{m>0}c^{(3)}_{m}q^{m}$ $=q^{3}-24q^{6}+O(q^{7}).$
Note that this basis is in reduced row echelon form in agreement with (19).
$\lambda_{2,1}=0,\quad\lambda_{2,2}=36,\quad\lambda_{2,3}=72$
for $Q_{4}$ and
$\lambda_{3,1}=2160,\quad\lambda_{3,2}=1728,\quad\lambda_{3,3}=1296,$
for $Q_{6}$ and their common eigenspaces $E_{1}$, $E_{2}$, and $E_{3}$ respectively.
5.5.3. Bounds on cusp forms in $\text{S}_{12}(\Gamma_{0}(3))$
We again use the estimates in [3] as collected in Section 4.2 to approximate the coefficients of $C^{(2)}_{12}(\tau)$ and $C^{(3)}_{12}(\tau)$.
The dimensions of $\mathcal{S}_{12}^{+}(\Gamma_{0}(3))$ and $\mathcal{S}_{12}^{-}(\Gamma_{0}(3))$ are given by
$d_{12}^{+}$ $=\left[\frac{12}{6}\right]=2,$
$d_{12}^{-}$ $=\left[\frac{12}{3}\right]-\left[\frac{12}{6}\right]-1=1.$
In principle this would require us to use Theorem 1.3 in [3] to obtain an upper bound on the coefficients of elements of $\mathcal{S}_{12}(\Gamma_{0}(\ell))$. However, in [3, Example 1.5.] this computation is done for arbitrary $C=\sum_{m=1}^{\infty}c_{m}q^{m}\in\mathcal{S}_{12}(\Gamma_{0}(3))$ and a bound on $|c_{m}|$ is given purely in terms of the first three coefficients $c_{1},c_{2},c_{3}$
$|c_{m}|$ $\leq d(m)m^{11/2}\left(e^{62.21}\left|\frac{82}{3^{4}}c_{1}+\frac{4}{3^{5}}c_{2}+\frac{1}{3^{6}}c_{3}\right|\right.$
$+\left.e^{39.39}\left|\frac{616}{3^{3}}c_{1}+\frac{130}{3^{4}}c_{2}-\frac{8}{3^{5}}c_{3}\right|+e^{54.08}\left|\frac{80}{3^{4}}c_{1}-\frac{4}{3^{5}}c_{2}-\frac{1}{3^{6}}c_{3}\right|\right).$
With $d(m)\leq 2\sqrt{m}$ we get
(52) $|c_{2,m}|\leq 3.5\cdot 10^{25}\cdot m^{6}\quad\text{and}\quad|c_{3,m}|\leq 2.9\cdot 10^{24}\cdot m^{6}$
for the coefficients of the cusp forms $C^{(2)}_{12}(\tau)$ and $C^{(3)}_{12}(\tau)$ respectively.
5.5.4. Eigenvalues of the Hessian of $L_{16}$
Explicitly, by Theorem 3.3, the eigenvalues of the Hessian are
(53) $\lambda_{j}=\frac{1}{288}\sum_{m=1}^{\infty}\left(\left(\sum_{i=2}^{3}\frac{9}{2}c_{i,m}\left(288\lambda_{i,j}-8i^{2}a_{i}\right)\right)+6a_{m}m(6m-9)\right)e^{-6m}$
Before using the previously found bounds for a rigorous verification of the fact that $L_{16}$ is a local maximum for the Gaussian potential given by $\alpha=3$, we plot the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,L_{16})$ at $L_{16}$. For this we use the first 200 summands of (24), the result is depicted in Figure 1.
Figure 1. Eigenvalues of the Hessian for $L_{16}$
From the plot we take numerical motivation to do the rigorous computation for $\alpha=3$ and the largest eigenvalue given by the orange curve, associated to the common eigenspace $E_{3}$ with the eigenvalues $\lambda_{2,3}=72$ and $\lambda_{3,3}=1296$ of the quadratic forms $Q_{4}$ and $Q_{6}$ respectively.
Now we give a rigorous argument that $\lambda_{3}<0$ for $\alpha=3$. For this we compute the sum of the first $50$ summands for $\lambda_{3}$ and obtain that this is bounded from above by $-0.2467$. We compute the tail of the series defining $\lambda_{3}$. By plugging in the explicit eigenvalues $\lambda_{2,3}=72$ and $\lambda_{3,3}=1296$ we find that the infinite sum in the tail is bounded by
$\sum_{m=51}^{\infty}\left(\tfrac{9}{2}\cdot 288(72|c_{2,m}|+1296|c_{3,m}|)+(6m)^{2}|a_{m}|\right)e^{-6m}$
$=$ $\sum_{m=51}^{\infty}\left(93312|c_{2,m}|+1679616|c_{3,m}|e^{-6m}+36|a_{m}|m^{2}\right)e^{-6m}$
[…]
where the inequality is obtained by using the estimates on the coefficients in (52) and (51). Applying Lemma 5.2 to the sums yields
$\sum_{m=51}^{\infty}m^{6}e^{-6m}\leq 4.6\cdot 10^{-123},$
$\sum_{m=51}^{\infty}m^{9}e^{-6m}\leq 6.1\cdot 10^{-118}.$
Putting everything together, we finally get that the tail is bounded by
$\frac{1}{288}(8.2\cdot 10^{30}\cdot 4.6\cdot 10^{-123}+1.2\cdot 10^{5}\cdot 6.1\cdot 10^{-118})$
$\leq$ $0.2467\displaystyle 1.3\cdot 10^{-94}<$ 0.2467 .
This concludes the proof that the eigenvalue of the Hessian of $\mathcal{E}(3,L_{16})$ associated to the eigenvalue pair $\lambda_{2,3}=72$ and $\lambda_{3,3}=1296$ of the forms $Q_{4}$ and $Q_{6}$, is negative. Similar computations for the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ for the common eigenspaces $E_{1}$ and $E_{2}$ show that $\lambda_{1}<0$ and $\lambda_{2}<0$, which makes the lattice $L_{16}$ a local maximizer for the energy function $L\mapsto\mathcal{E}(3,L)$. ∎

### v2

@@ 5.5.1. Bounding the coefficients of $\Theta_{L_{16}}$
$|b^{(1)}|,|b^{(2)}_{m}|\leq|b_{m}|\leq 485m^{7}$
Again, $\mathcal{S}_{8}(\Gamma_{0}(3))$ is $1$-dimensional, so $C_{8}(\tau)=\sum_{m=1}^{\infty}c_{m}q^{m}$ is a Hecke eigenform, and (35) gives
$|c_{m}|$ $\leq 2m^{4}.$
(39) $|a_{m}|\leq 3325\cdot m^{7}+12\cdot m^{4}.$
5.5.2. The auxiliary forms $Q_{4}$ and $Q_{6}$
$Q_{2m}=\sum_{x\in L_{16}(2m)}H[x]^{2}$
specifically for $m=1,2,3$.
$C^{(3)}_{12}(\tau)=\sum_{m>0}c^{(3)}_{m}q^{m}$ $=q^{3}-24q^{6}+O(q^{7}).$
Note that this basis is in reduced row echelon form in agreement with (19). However, this is not a basis of Hecke eigenforms. One such basis can be obtained using the mfbasis command in PARI/GP [26] we compute the following basis of Hecke eigenforms
[…]
These are in fact Hecke eigenforms. Firstly $H^{(1)}_{12}$ and $H^{(2)}_{12}$ are expressed in terms of Ramanujan’s discriminant function $\Delta$, while the third one is collected in the L-functions and modular forms database [20], as entry Newform orbit 3.12.a.a.
$\lambda_{2,1}=0,\quad\lambda_{2,2}=72,\quad\lambda_{2,3}=144$
for $Q_{4}$ and
$\lambda_{3,1}=4320,\quad\lambda_{3,2}=3456,\quad\lambda_{3,3}=2592,$
for $Q_{6}$ and their common eigenspaces $E_{1}$, $E_{2}$, and $E_{3}$ respectively.
For the coefficients of $C^{(2)}_{12}$ and $C^{(3)}_{12}$ we express these cusp forms in the basis (40) of Hecke eigenforms:
$C^{(2)}_{12}=\frac{1}{102}(-H^{(1)}_{12}+165H^{(2)}_{12}+H^{(3)}_{12})$
and
$C^{(3)}_{12}=H^{(2)}_{12}$.
Using (35) for the coefficients of the Hecke eigenforms we find
(41) $|c_{2,m}|\leq(1+165+1)\cdot 2\cdot m^{6}=334\cdot m^{6}\quad\text{and}\quad|c_{3,m}|\leq 2\cdot m^{6}$
for the coefficients of the cusp forms $C^{(2)}_{12}(\tau)$ and $C^{(3)}_{12}(\tau)$ respectively. Note that here we could strengthen the bound on $C^{{3}}_{12}$, as $C^{{3}}_{12}(\tau)=H^{(2)}_{12}=\Delta(3\tau)$, where the latter is a Hecke eigenform for $\mathcal{S}_{12}(\operatorname{SL}_{2}(\mathbb{Z}))$, so its coefficients are bound by $d(m/3)\cdot(m/3)^{11/2}\leq 2\cdot 3^{-6}m^{6}$, rather than by $2\cdot m^{6}$. However, we will not make use of this strengthened version fur the current application.
5.5.3. Eigenvalues of the Hessian of $L_{16}$
Explicitly, by Theorem 3.3, the eigenvalues of the Hessian are
(42) $\lambda_{j}=\frac{1}{288}\sum_{m=1}^{\infty}\left(\left(\sum_{i=2}^{3}\frac{1}{2}c_{i,m}\left(288\lambda_{i,j}-8i^{2}a_{i}\right)\right)+2a_{m}m(2m-9)\right)e^{-2m}$
Before using the previously found bounds for a rigorous verification of the fact that $L_{16}$ is a local maximum for the Gaussian potential given by $\alpha=1$, we plot the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,L_{16})$ at $L_{16}$. For this we use the first 200 summands of (24), the result is depicted in Figure 1.
Figure 1. Eigenvalues of the Hessian for $L_{16}$
We provide a rigorous estimate at $\alpha=1$ to confirm the numerical simulation. Specifically, we provide two estimates. First, we show that the numerically negative eigenvalue $\lambda_{1}$, which is associated to the common eigenspace $E_{1}$ with the eigenvalues $\lambda_{2,1}=0$ and $\lambda_{3,1}=4320$, and printed as the violet curve, is indeed strictly negative. Second, we show that the numerically smaller positive eigenvalue $\lambda_{2}$, which is associated to the common eigenspace $E_{2}$ with the eigenvalues $\lambda_{2,3}=72$ and $\lambda_{3,3}=2592$, and printed as the orange curve, is indeed strictly positive.
We first compute the sum of the first $20$ summands as an approximation for $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$:
$\lambda_{1}\approx-0.062,\quad\lambda_{2}\approx 0.024,\quad\text{and}\quad\lambda_{3}\approx 0.111.$
To provide a rigorous proof of negativity and positivity we now compute bounds on the tails $T_{1}$, $T_{2}$, and $T_{3}$, of the series defining $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ respectively. These bounds will show that the tail is only affecting the eigenvalue to the order of $10^{-5}$.
For $\lambda_{1}$, we use the explicit eigenvalues $\lambda_{2,1}=0$ and $\lambda_{3,1}=4320$, and obtain that
$|T_{1}|\leq$ $\sum_{m=21}^{\infty}\left(\tfrac{1}{2}\cdot 288(0\cdot|c_{2,m}|+4320\cdot|c_{3,m}|)+(2m)^{2}|a_{m}|\right)e^{-2m}$
$=$ $\sum_{m=21}^{\infty}\left(622080\cdot|c_{3,m}|+4\cdot|a_{m}|m^{2}\right)e^{-2m}$
$\leq\$$1.3\cdot 10^{6}\sum_{m=21}^{\infty}m^{6}e^{-2m}+1.4\cdot 10^{4}\sum_{m=21}^{\infty}m^{9}e^{-2m}$
where the inequality is obtained by using the estimates on the coefficients in (41) and (39). Similarly we obtain
$|T_{2}|\leq$ $\sum_{m=21}^{\infty}\left(\tfrac{1}{2}\cdot 288(72\cdot|c_{2,m}|+3456\cdot|c_{3,m}|)+(2m)^{2}|a_{m}|\right)e^{-2m}$
$\leq\$$4.5\cdot 10^{6}\sum_{m=21}^{\infty}m^{6}e^{-2m}+1.4\cdot 10^{4}\sum_{m=21}^{\infty}m^{9}e^{-2m}$
for $\lambda_{2}$ with $\lambda_{2,2}=72$ and $\lambda_{3,2}=3456$, and
$|T_{3}|\leq$ $\sum_{m=21}^{\infty}\left(\tfrac{1}{2}\cdot 288(144\cdot|c_{2,m}|+2592\cdot|c_{3,m}|)+(2m)^{2}|a_{m}|\right)e^{-2m}$
$\leq\$$7.7\cdot 10^{6}\sum_{m=21}^{\infty}m^{6}e^{-2m}+1.4\cdot 10^{4}\sum_{m=21}^{\infty}m^{9}e^{-2m}$
for $\lambda_{2}$ with $\lambda_{2,3}=144$ and $\lambda_{3,3}=2592$.
To conclude the estimation on the tails, we use the following lemma (Lemma 2.2 in [16]), which provides an analytic expression for the remaining infinite series.
Lemma 5.5.
For $j\geq\frac{k}{2\alpha}$, we have
where
$\Gamma(s,x)=\int_{x}^{\infty}t^{s-1}e^{-t}\,dt$
is the incomplete gamma function.
Applying Lemma 5.5 to the series in $T_{1}$, $T_{2}$, and $T_{3}$ yields
$\sum_{m=21}^{\infty}m^{6}e^{-2m}\leq 7.8\cdot 10^{-11},$
$\sum_{m=21}^{\infty}m^{9}e^{-2m}\leq 7.5\cdot 10^{-7}.$
With this we obtain the final bounds
$|T_{1}|\leq$ $\frac{1}{288}(1.3\cdot 10^{6}\cdot 7.8\cdot 10^{-11}+1.4\cdot 10^{4}\cdot 7.5\cdot 10^{-7})$
$\leq\$$3.7\cdot 10^{-5}<0.062,$
and analogously
$|T_{2}|\leq 3.8\cdot 10^{-5}<0.024\quad\text{and}\quad|T_{3}|\leq 3.9\cdot 10^{-5}<0.111.$
This concludes the proof. ∎
5.6. Extremal modular lattices which are not universally critical
@@
Note that, up to isometry, there are two more such lattices. For the first such lattice, the first entry in “2_dim12.dat” in [25] all shells form spherical $3$-designs; to test this one can use the design test based on harmonic Molien series as in Goethals and Seidel [15], which we explained in Section 2.4. So this lattice is a universal critical point.
@@
This follows the same lines as the computation in Section 5.5. Note that $\Theta_{L_{12}}$ is a modular form of weight 6 for $\Gamma_{0}(2)$. The space $\mathcal{M}_{6}(\Gamma_{0}(2))$ contains no cusp forms and is therefore equal to its two-dimensional subspace of Eisenstein series $\operatorname{Eis}_{6}(\Gamma_{0}(2))$.
@@
5.7. Numerical considerations
While we only discussed a number of examples in detail, we performed a numerical investigation of all extremal $2$-modular lattices up to dimension $20$ and $3$-modular lattices up to dimension $18$. We briefly collect the qualitative behavior of these lattices (in the space of lattices of fixed density) in the Gaussian core model, the techniques used in the preceding examples can be used to rigorously verify these observations, if necessary.
The relevant data is presented in Tables 4 and 5 for $2$ and $3$-modular lattices separately. We collect for each dimension how such lattices behave in the Gaussian core model (in the space of lattices of fixed density) according to our numerical computations, and give a reference to a (partial) proof of the claims, if it is available. If there are multiple isometry classes of such lattices in a given dimension it is usually the case that the number of distinct eigenvalues of the Hessian depends on the lattice.
In some cases, $\ell$-modular lattices of the same dimension behave differently: We observed this in dimension $12$ for $2$-modular lattices, and in dimensions $8$ and $10$ for $3$-modular lattices. Here, the minimal design strength of the shells, as recorded in Tables 1 and 2, is not big enough to guarantee that we have at least $2$-designs on every shell. We therefore investigated the design strength in a case by case study, by computing the harmonic Molien series in every case, and using Theorem 3.12 in [15], as explained in Section 2.4. It turns out, that in all three cases, precisely one lattice is universally critical, the corresponding shells are always at least $3$-designs, while the rest are not universally critical, the design strength of some shells is $1$. In the table we use the notation $d.i$ to refer to the $i$th lattice in dimension $d$ as given in the database [25].
dimension lattices behavior reference
44 all loc. univ. opt. [30], Theorem 5.3
88 all univ. saddle point
$12$ 12.1 univ. saddle point
12.2, 12.3 not univ. critical Proposition 5.6
$16$ all loc. univ. opt. Proposition 5.2
2020 all loc. univ. opt. Theorem 5.3
Table 4. Numerical data on the behavior of $2$-modular extremal lattices in small dimensions.
dimension lattices behavior reference
22 all (loc.) univ. opt. [23], Theorem 5.3
44 all univ. saddle point
66 all univ. saddle point
88 8.1 univ. saddle point
8.2 not univ. critical
$10$ 10.1 univ. saddle point
10.2, 10.3 not univ. critical
$12$ all loc. univ. opt. Proposition 5.1
$14$ all loc. univ. opt. Theorem 5.3
$16$ all univ. saddle point
$18$ all univ. saddle point
Table 5. Numerical data on the behavior of $3$-modular extremal lattices in small dimensions.